Calculus II, Fall 2014 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 03

1. Section 10.8: Problems 7, 9, 29, 33, 35, 41.

2. Let

f(x) = \{ \begin{matrix}0, & x = 0 \\ e^{−1/x^{2}}, & x ≠ 0 \end{matrix}

it is known that f^{(n)}(0) = 0 for all n. Verify this for f′(0) and f′′(0).

3. Section 10.9: Problems 7, 9, 17, 19, 33, 41, 42, 50(a).